OpenStudy (anonymous):

It takes 242.kj/mol to break a chlorine-chlorine single bond. Calculate the maximum wavelength of light for which a chlorine-chlorine single bond could be broken by absorbing a single photon.

OpenStudy (nincompoop):

CONVERT kj per mol into j per bond, which is equivalent to j per photon. This is going to be your E \[\frac{242. kJ}{mol} \times \frac{1mol}{6.022 \times 10 ^{23} bonds} \times \frac{1000 mol}{1kJ}\] THEN use the energy per photon formula to calculate for the wavelength \[E=\frac{hc}{\lambda } \rightarrow \lambda = \frac{hc}{E}\] where: \[h = 6.626 \times 10^{34} \frac{J}{s}\] \[c \approx 3.0 \times 10^{8} \frac{m}{s}\]

OpenStudy (jfraser):

@nincompoop you are missing a negative sign in your exponent for planck's constant, it should be \[h = 6.626*10^{-34}J*s\]

OpenStudy (nincompoop):

ooops thanks